OpenStudy (anonymous):

Help me simplify this. 8x^2+4x+2/1-8^2

OpenStudy (anonymous):

That can't really be simplified to nicely.

OpenStudy (mertsj):

\[\frac{8x^2+4x+2}{1-8^2}=\frac{9x^2+4x+2}{-63}\]

OpenStudy (mertsj):

Whoops. 8x^2 in numerator.

OpenStudy (anonymous):

use this formula to factor the denominator A^3-B^3=(A-B)(A^2+AB+B^2)

OpenStudy (anonymous):

it's 8X^3 not 8x^ cubed not squared

OpenStudy (mertsj):

why don't you just start over and post the problem correctly?

OpenStudy (anonymous):

I only noticed it when you typed it sorry

OpenStudy (anonymous):

It's just cubed for the denominator, the numerator is squred

OpenStudy (mertsj):

\[\frac{8x^2+4x+2}{1-8x^3}\]

OpenStudy (mertsj):

Is that the correct problem?

OpenStudy (anonymous):

yes, it is the correct proble. Can you do it?

OpenStudy (mertsj):

\[\frac{8x^2+4x+2}{1-8x^3}=\frac{2(4x^2+2x+1)}{(1-2x)(1+2x+4x^2)}=\frac{2}{1-2x}\]

OpenStudy (anonymous):

yur a genius! u get a medal! thanhank u :)

OpenStudy (mertsj):

yw.

OpenStudy (anonymous):

what does yw mean?

OpenStudy (mertsj):

You're welcome

OpenStudy (anonymous):

oh :)